OpenStudy (anonymous):

1. Which of the following atoms would you expect to release less energy than carbon when an electron is added to its valence? oxygen (O) fluorine (F) chlorine (Cl) neon (Ne) 2. Consider the compound Al(NO3)3 and answer the following questions. What is the name of this compound? What is the molar mass of this compound? Explain how you calculated this value. 3. Suppose a laboratory wants to identify an unknown pure substance. The valence electrons of the substance's atoms feel an effective nuclear charge of +4. If the unknown substance has a higher electronegativity value than silicon (Si), what other characteristics would you expect the substance to exhibit? It would have larger atomic radii than Si and higher ionization energies than Si. It would have larger atomic radii than Si and lower ionization energies than Si. It would have smaller atomic radii than Si and higher ionization energies than Si. It would have smaller atomic radii than Si and lower ionization energies than Si. 4. The trend for ionization energy is a general increase from left to right across a period. However, nitrogen (N) is found to have a higher first ionization energy value than oxygen (O). Explain this exception to the general trend in terms of electron arrangements and attraction/repulsion. 5. A hydrate of Na2CO3 has a mass of 4.31 g before heating. After heating, the mass of the anhydrous compound is found to be 3.22 g. Determine the formula of the hydrate and then write out the name of the hydrate. Please show all your work for the calculations for full credit.

OpenStudy (unklerhaukus):

i can help with the second part of 2

OpenStudy (anonymous):

okay

OpenStudy (unklerhaukus):

The molar mass of \(\text{Al}(\text{NO}_3)_3\) \[M_{\text{Al}(\text{NO}_3)_3}=M_\text{Al}+3M_{\text{NO}_3}\\ \qquad\qquad=M_\text{Al}+3(M_{\text N}+3M_\text O)\\ \qquad\qquad=M_\text{Al}+3M_{\text N}+9M_\text O\]

OpenStudy (anonymous):

i dont understand that

OpenStudy (unklerhaukus):

do you know what molar mass means?

OpenStudy (anonymous):

yea but like i dont understand what you just wrote

OpenStudy (unklerhaukus):

I use \(M_\text x\) to denote the molar mass of a chemical \(\text x\)

OpenStudy (unklerhaukus):

The molar masses of elements \(M_x\) can be found on a periodic table I broke the chemical into its elements to make the calculation simpler

OpenStudy (anonymous):

well what would the answer be?

OpenStudy (unklerhaukus):

on my periodic table \[M_\text{Al}=27\left[\tfrac{\text g}{\text{mol}}\right]\\M_{\text N}=14\left[\tfrac{\text g}{\text{mol}}\right]\\M_\text O=16\left[\tfrac{\text g}{\text{mol}}\right]\]

OpenStudy (unklerhaukus):

so 27 + 3x14 + 9x16

OpenStudy (anonymous):

is that 3 times 14 and 9 times 16 or are they x's

OpenStudy (unklerhaukus):

\times

OpenStudy (anonymous):

okay so that would be 27 + 42 + 144 or 213

OpenStudy (unklerhaukus):

yeah

OpenStudy (anonymous):

so the molar mass is 213?

OpenStudy (unklerhaukus):

yes

OpenStudy (anonymous):

okay uhm can you help me with any of the others?